# RaSeRec: Retrieval-Augmented Sequential Recommendation

### v2

@@ 3.1 Collaborative-based Pre-training
Retrieval Training. To retrieve memories that share the same (or similar) preference as the input user sequence, a retrieval training objective is developed to maximize the agreement of positive pairs and minimize that of negative pairs. The positive pair consists of two user sequences that share the same (or similar) preference. In SeRec, the goal is to predict the next item. Therefore, if the next item of two user sequences is the same, these two sequences can be thought to have similar collaborative semantics Qiu et al. (2022). Particularly, the positive pair can be defined as $s_{u^{\prime},t^{i}}=[v_{t^{i}-T+1},...,v_{t^{i}}],$ and $s_{u^{\prime\prime},t^{j}}=[v_{t^{j}-T+1},...,v_{t^{j}}]$, where $v_{t^{i}+1}$ and $v_{t^{j}+1}$ are the same item. For brevity, we mark $s_{u^{\prime},t^{i}},s_{u^{\prime\prime},t^{j}}$ as $s^{\prime},s^{\prime\prime}$, respectively. The user representation of them can be defined as:
$\mathbf{h}^{\prime}=\mathrm{SeqEnc(s^{\prime})},\quad\mathbf{h}^{\prime\prime}=\mathrm{SeqEnc(s^{\prime\prime})}.$ (6)
@@
In retrieval training settings, negative samples usually need to be selected from a large pool Karpukhin et al. (2020). To efficiently construct negatives, we treat the other samples in the same training batch as negatives. Assume that we have a training batch $\mathcal{B}$ that consists of $|\mathcal{B}|$ positive pairs:
=$\mathcal{B}=\{\langle\mathbf{h}^{\prime}_{1},\mathbf{h}^{\prime\prime}_{1}\rangle,\langle\mathbf{h}^{\prime}_{2},\mathbf{h}^{\prime\prime}_{2}\rangle,...,\langle\mathbf{h}^{\prime}_{|\mathcal{B}|},\mathbf{h}^{\prime\prime}_{|\mathcal{B}|}\rangle\}.$ (7)
@@
As such, for each positive pair, there are $2(|\mathcal{B}|-1)$ negatives forming the negative set $\mathcal{S}^{-}$. For example, for the positive pair $\langle\mathbf{h}^{\prime}_{1},\mathbf{h}^{\prime\prime}_{1}\rangle$, its corresponding negative set is $\mathcal{S}^{-}_{1}=\{\mathbf{h}^{\prime}_{2},\mathbf{h}^{\prime\prime}_{2},...,\mathbf{h}^{\prime}_{|\mathcal{B}|},\mathbf{h}^{\prime\prime}_{|\mathcal{B}|}\}$. Having established the positive and negative pairs, we adopt InfoNCE Gutmann and Hyvärinen (2010) as the loss function, which can be defined as follows:
$\begin{split}\mathcal{L}_{ret}=&-\Big{(}{\log\frac{e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{\prime\prime}_{i})/\tau}}{e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{\prime\prime}_{i})/\tau}+\sum_{s^{-}\in\mathcal{S}^{-}_{i}}e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{-})/\tau}}}\\ &+{\log\frac{e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{\prime}_{i})/\tau}}{e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{\prime}_{i})/\tau}+\sum_{s^{-}\in\mathcal{S}^{-}_{i}}e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{-})/\tau}}}\Big{)},\end{split}$ (8)
@@
Figure 3: The architecture of Retrieval-Augmented Module, where $\hat{\mathbf{h}}$ is the augmented user representation.
3.2 Retrieval-Augmented Fine-tuning
@@
Memory Retrieval. Collaborative <user sequence $s_{u,t}$, target item $v_{t+1}^{(u)}$> pairs explicitly reveal the sequential patterns, where pairs sharing the same (or similar) preference as the input user sequence can act as important references. Inspired by the above idea, we construct a reference set $\mathcal{R}$ by auto-regressively enumerating all <$s_{u,t}$, $v_{t+1}^{(u)}$> pairs from the training data except the current input user sequence. We then employ the sequence encoder $\mathrm{SeqEnc}(\cdot)$ to encode all pairs in $\mathcal{R}$ so that we can get a memory bank $\mathcal{M}$, where each entry in $\mathcal{M}$ consists of a user representation and the corresponding target item embedding. Given the input representation $\mathbf{h}$, RaSeRec retrieves top-$K$ similar user representations $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ as well as their corresponding target item embedding $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ from the memory bank $\mathcal{M}$, where we employ the Faiss library Johnson et al. (2021) to speed up the  retrieval  process.
Representation Augmentation. With the retrieved memories, we design a Retrieval-Augmented Module (RAM), which learns to leverage the retrieved memories to augment the current user representation. Its overall architecture is illustrated in Figure 3. Inspired by Xie et al. (2023), RAM employs dual-channel multi-head cross attention (MHCA) which takes $\mathbf{h}$ as query, $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ (or $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$) as key and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ (or $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$) value to get the augmented user representation. In the first channel, RAM learns to weight and aggregate $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ via modeling the relationship between $\mathbf{h}$ and $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$:
$\mathbf{h}^{c_{1}}=\mathrm{MHCA}(\mathbf{h},\{{\mathbf{h}}_{k}\}_{k=1}^{K},\{{\mathbf{v}}_{k}\}_{k=1}^{K}),$ (9)
@@
$\mathbf{h}^{c_{2}}=\mathrm{MHCA}(\mathbf{h},\{{\mathbf{v}}_{k}\}_{k=1}^{K},\{{\mathbf{h}}_{k}\}_{k=1}^{K}).$ (10)
@@
$\tilde{\mathbf{h}}=\alpha\mathbf{h}+(1-\alpha)(\beta\mathbf{h}^{c_{1}}+(1-\beta)\mathbf{h}^{c_{2}}),$ (11)
where $0\leq\alpha,\beta\leq 1$ control the strength of each representation. A possible extension is to learn $\alpha$ and $\beta$, e.g., designing attention mechanism Wu et al. (2019). However, it is not the focus of this work, and we leave it for future exploration. We further fine-tune the model with RAFT to improve recommendations with the retrieval-augmented representation $\tilde{\mathbf{h}}$. Specifically, we freeze the model backbone $\mathrm{SeqEnc}(\cdot)$ and only the RAM is being updated during fine-tuning. Formally, we adopt the same objective as Equ (5), which learns to leverage the retrieved memories for better recommendation:
$\mathcal{L}_{raft}=-\log\Big{(}e^{\tilde{\mathbf{h}}^{T}\mathbf{v}_{t+1}}/\sum_{v_{i}\in\mathcal{V}}e^{\tilde{\mathbf{h}}^{T}\mathbf{v}_{i}}\Big{)}.$ (12)
@@
Table 1: The space and time complexity comparison between the Vanilla and SSL/Retrieval-Augmented SeRec.
3.3 Inference and Complexity Analyses
@@
In this section, we first introduce the model inference process and then analyze the model complexity in terms of both time as well as space in detail.
@@
Model Inference. During inference, RaSeRec first computes the user representation $\mathbf{h}$ via $\mathrm{SeqEnc}(\cdot)$ for the input user sequence $s_{u}$. Then, RaSeRec retrieves top-$K$ similar memories (i.e., $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$) from the memory bank $\mathcal{M}$. After that, RaSeRec employs RAM that consists of dual-channel MHCA to weight and aggregate valuable patterns from both $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$. We obtain the final retrieval-augmented user representation $\tilde{\mathbf{h}}$ by combining both the original user representation $\mathbf{h}$ and the augmented user ones of the dual channels i.e., $\mathbf{h}^{c_{1}}$ and $\mathbf{h}^{c_{2}}$. Lastly, we recommend the top-$N$ items based on the dot product scores between $\mathbf{h}$ and item  embedding  matrix $\mathbf{V}$.
@@
Complexity Analyses. Table 1 summarizes the space and time complexity of three SeRec paradigms, where we compare the complexity of RaSeRec with the Vanilla and SSL-Augmented SeRec to manifest the efficiency of RaSeRec. For space complexity, RaSeRec needs only $O(|\phi|+|\theta|)$, where $\phi$ and $\theta$ represent the trainable parameters of the RAM and sequence encoder, respectively. Generally, $|\phi|$ is far below than $|\theta|$. As such, the analytical space complexity of RaSeRec is in the same magnitude as the Vanilla and SSL-Augmented one ($O(|\theta|)$). In practice, taking the Amazon Sports dataset as an example, the space occupation of RaSeRec is merely  0.05  times  larger  than  theirs.
@@
As recommender systems mainly focus on the time cost of online serving, we analyze the time complexity during inference. The time complexity for encoding user representations is $O(L(Td^{2}+T^{2}d))$, where $T$ is the maximum sequence length and $d$ is the hidden size. In addition to encoding, RaSeRec also needs to retrieve memories and augment user representation. For retrieval, RaSeRec instantiates Faiss using Inverted File Indexing (IVF), whose retrieval complexity is $O(kd+\frac{|\mathcal{V}|}{k}d)$, where $k$ denotes the number of clusters and we set the number of clusters accessed as 1. On the other hand, the time complexity of augmenting user representation is $O(Kd)$. Finally, these three SeRec paradigms need to compute dot product scores over all items to make recommendations, resulting in a time complexity $O(|\mathcal{V}|d)$. The time complexity of additional retrieval and augmentation processes (i.e., $O(kd+\frac{|\mathcal{V}|}{k}d+Kd)$) is much lower than the encoding and recommendation processes (i.e., $O(L(Td^{2}+T^{2}d)+|\mathcal{V}|d)$). Specifically, $k+\frac{|\mathcal{V}|}{k}+K$ is on the order of thousands, while $|\mathcal{V}|$ is on the order of tens of thousands. Therefore, the analytical time complexity of RaSeRec is in the same magnitude as the Vanilla  and  SSL-Augmented  SeRec.
@@
Table 2: Performance comparison (@5), where the best results are boldfaced and the second-best ones are underlined.
@@
Table 3: Performance comparison in terms of different sequential recommendation model backbones.
4 Experiments
To verify the effectiveness of RaSeRec, we conduct extensive experiments on three benchmark datasets to answer the following research questions (RQs): RQ1: How does RaSeRec perform compared with state-of-the-art SeRec models? RQ2: How much gain can RaSeRec bring to the existing base backbones? RQ3: Can RaSeRec improve long-tailed recommendation? More in-depth studies about RaSeRec can be found in the Appendices.
4.1 Experimental Settings
@@
Datasets. We conduct experiments on three benchmark datasets collected from Amazon McAuley et al. (2015). We adopt three sub-categories, i.e., Beauty, Sports, and Clothing [^2] http://jmcauley.ucsd.edu/data/amazon/, with different sparsity degrees. Following Kang and McAuley (2018); Sun et al. (2019), we discard users and items with fewer than 5 interactions. The statistics of datasets after preprocessing  are  provided  in  Appendix B.

### v3

@@ 3.1.2. Retrieval Training
To retrieve memories that share the same (or similar) preference as the input user sequence, a retrieval training objective is developed to maximize the agreement of positive pairs and minimize that of negative pairs. The positive pair consists of two user sequences that share the same (or similar) preference. In SeRec, the goal is to predict the next item. Therefore, if the next item of two user sequences is the same, these two sequences can be thought to have similar collaborative semantics (Qiu et al., 2022). Particularly, the positive pair can be defined as $s_{u^{\prime},t^{i}}=[v_{t^{i}-T+1},...,v_{t^{i}}],$ and $s_{u^{\prime\prime},t^{j}}=[v_{t^{j}-T+1},...,v_{t^{j}}]$, where $v_{t^{i}+1}$ and $v_{t^{j}+1}$ are the same item. For brevity, we mark $s_{u^{\prime},t^{i}},s_{u^{\prime\prime},t^{j}}$ as $s^{\prime},s^{\prime\prime}$, respectively. The user representation of them can be defined as:
(6) $\mathbf{h}^{\prime}=\mathrm{SeqEnc(s^{\prime})},\quad\mathbf{h}^{\prime\prime}=\mathrm{SeqEnc(s^{\prime\prime})}.$
@@
In retrieval training settings, negative samples usually need to be selected from a large pool (Karpukhin et al., 2020). To efficiently construct negatives, we treat the other samples in the same training batch as negatives. Assume that we have a training batch $\mathcal{B}$ that consists of $|\mathcal{B}|$ positive pairs:
(7) =$\mathcal{B}=\{\langle\mathbf{h}^{\prime}_{1},\mathbf{h}^{\prime\prime}_{1}\rangle,\langle\mathbf{h}^{\prime}_{2},\mathbf{h}^{\prime\prime}_{2}\rangle,...,\langle\mathbf{h}^{\prime}_{|\mathcal{B}|},\mathbf{h}^{\prime\prime}_{|\mathcal{B}|}\rangle\}.$
@@
As such, for each positive pair, there are $2(|\mathcal{B}|-1)$ negatives forming the negative set $\mathcal{S}^{-}$. For example, for the positive pair $\langle\mathbf{h}^{\prime}_{1},\mathbf{h}^{\prime\prime}_{1}\rangle$, its corresponding negative set is $\mathcal{S}^{-}_{1}=\{\mathbf{h}^{\prime}_{2},\mathbf{h}^{\prime\prime}_{2},...,\mathbf{h}^{\prime}_{|\mathcal{B}|},\mathbf{h}^{\prime\prime}_{|\mathcal{B}|}\}$. Having established the positive and negative pairs, we adopt InfoNCE (Gutmann and Hyvärinen, 2010) as the loss function, which can be defined as follows:
(8) $\begin{split}\mathcal{L}_{ret}=&-\Big{(}{\log\frac{e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{\prime\prime}_{i})/\tau}}{e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{\prime\prime}_{i})/\tau}+\sum_{s^{-}\in\mathcal{S}^{-}_{i}}e^{\mathrm{s}(\mathbf{h}^{\prime}_{i},\mathbf{h}^{-})/\tau}}}\\ &+{\log\frac{e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{\prime}_{i})/\tau}}{e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{\prime}_{i})/\tau}+\sum_{s^{-}\in\mathcal{S}^{-}_{i}}e^{\mathrm{s}(\mathbf{h}^{\prime\prime}_{i},\mathbf{h}^{-})/\tau}}}\Big{)},\end{split}$
@@
Figure 3. The architecture of Retrieval-Augmented Module, where $\tilde{\mathbf{h}}$ is the augmented user representation.
3.2. Retrieval-Augmented Fine-tuning
@@
3.2.1. Memory Retrieval
Collaborative ¡user sequence $s_{u,t}$, target item $v_{t+1}^{(u)}$¿ pairs explicitly reveal the sequential patterns, where pairs sharing the same (or similar) preference as the input user sequence can act as important references. Inspired by the above idea, we construct a reference set $\mathcal{R}$ by auto-regressively enumerating all ¡$s_{u,t}$, $v_{t+1}^{(u)}$¿ pairs from the training data except the current input user sequence. We then employ the sequence encoder $\mathrm{SeqEnc}(\cdot)$ to encode all pairs in $\mathcal{R}$ so that we can get a memory bank $\mathcal{M}$, where each entry in $\mathcal{M}$ consists of a user representation and the corresponding target item embedding. Given the input representation $\mathbf{h}$, RaSeRec retrieves top-$K$ similar user representations $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ as well as their corresponding target item embedding $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ from the memory bank $\mathcal{M}$, where we employ the Faiss library (Johnson et al., 2021) to speed up the retrieval process.
3.2.2. Representation Augmentation
With the retrieved memories, we design a Retrieval-Augmented Module (RAM), which learns to leverage the retrieved memories to augment the current user representation. Its overall architecture is illustrated in Figure 3. Inspired by (Xie et al., 2023), RAM employs dual-channel multi-head cross attention (MHCA) which takes $\mathbf{h}$ as query, $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ (or $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$) as key and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ (or $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$) as value to get the augmented user representation. In the first channel, RAM learns to weight and aggregate $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$ via modeling the relationship between $\mathbf{h}$ and $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$:
(9) $\mathbf{h}^{c_{1}}=\mathrm{MHCA}(\mathbf{h},\{{\mathbf{h}}_{k}\}_{k=1}^{K},\{{\mathbf{v}}_{k}\}_{k=1}^{K}),$
@@ 3.2.2. Representation Augmentation
(10) $\mathbf{h}^{c_{2}}=\mathrm{MHCA}(\mathbf{h},\{{\mathbf{v}}_{k}\}_{k=1}^{K},\{{\mathbf{h}}_{k}\}_{k=1}^{K}).$
@@
(11) $\tilde{\mathbf{h}}=\alpha\mathbf{h}+(1-\alpha)(\beta\mathbf{h}^{c_{1}}+(1-\beta)\mathbf{h}^{c_{2}}),$
where $0\leq\alpha,\beta\leq 1$ control the strength of each representation. A possible extension is to learn $\alpha$ and $\beta$, e.g., designing attention mechanism (Wu et al., 2019). However, it is not the focus of this work, and we leave it for future exploration. We further fine-tune the model with RAFT to improve recommendations with the retrieval-augmented representation $\tilde{\mathbf{h}}$. Specifically, we freeze the model backbone $\mathrm{SeqEnc}(\cdot)$ and only the RAM is being updated during fine-tuning. Formally, we adopt the same objective as Equ (5), which learns to leverage the retrieved memories for better recommendation:
(12) $\mathcal{L}_{raft}=-\log\Big{(}e^{\tilde{\mathbf{h}}^{T}\mathbf{v}_{t+1}}/\sum_{v_{i}\in\mathcal{V}}e^{\tilde{\mathbf{h}}^{T}\mathbf{v}_{i}}\Big{)}.$
In this way, RaSeRec does not have to memorize all long-tailed patterns but learns to leverage retrieved explicit memories to accommodate preference drift.
Table 1. The space and time complexity comparison between the Vanilla and SSL/Retrieval-Augmented SeRec.
@@
3.3. Inference and Complexity Analyses
@@
In this section, we first introduce the model inference process and then analyze the model complexity in terms of both time and space in detail.
3.3.1. Model Inference
@@
During inference, RaSeRec first computes the user representation $\mathbf{h}$ via $\mathrm{SeqEnc}(\cdot)$ for the input user sequence $s_{u}$. Then, RaSeRec retrieves top-$K$ similar memories (i.e., $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$) from the memory bank $\mathcal{M}$. After that, RaSeRec employs RAM that consists of dual-channel MHCA to weight and aggregate valuable patterns from both $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$. We obtain the final retrieval-augmented user representation $\tilde{\mathbf{h}}$ by combining both the original user representation $\mathbf{h}$ and the augmented user ones of the dual channels i.e., $\mathbf{h}^{c_{1}}$ and $\mathbf{h}^{c_{2}}$. Lastly, we recommend the top-$N$ items based on the dot product scores between $\tilde{\mathbf{h}}$ and item embedding matrix $\mathbf{V}$.
3.3.2. Complexity Analyses
@@
Table 1 summarizes the space and time complexity of three SeRec paradigms, where we compare the complexity of RaSeRec with the Vanilla and SSL-Augmented SeRec to manifest the efficiency of RaSeRec. For space complexity, RaSeRec needs only $O(|\phi|+|\theta|)$, where $\phi$ and $\theta$ represent the trainable parameters of the RAM and sequence encoder, respectively. Generally, $|\phi|$ is far below than $|\theta|$. As such, the analytical space complexity of RaSeRec is in the same magnitude as the Vanilla and SSL-Augmented one ($O(|\theta|)$). In practice, taking the Amazon Sports dataset as an example, the space occupation of RaSeRec is merely 0.05 times larger than theirs.
@@
As recommender systems mainly focus on the time cost of online serving, we analyze the time complexity during inference. The time complexity for encoding user representations is $O(L(Td^{2}+T^{2}d))$, where $T$ is the maximum sequence length and $d$ is the hidden size. In addition to encoding, RaSeRec also needs to retrieve memories and augment user representation. For retrieval, RaSeRec instantiates Faiss using Inverted File Indexing (IVF), whose retrieval complexity is $O(kd+\frac{|\mathcal{V}|}{k}d)$, where $k$ denotes the number of clusters and we set the number of clusters accessed as 1. On the other hand, the time complexity of augmenting user representation is $O(Kd)$. Finally, these three SeRec paradigms need to compute dot product scores over all items to make recommendations, resulting in a time complexity $O(|\mathcal{V}|d)$. The time complexity of additional retrieval and augmentation processes (i.e., $O(kd+\frac{|\mathcal{V}|}{k}d+Kd)$) is much lower than the encoding and recommendation processes (i.e., $O(L(Td^{2}+T^{2}d)+|\mathcal{V}|d)$). Specifically, $k+\frac{|\mathcal{V}|}{k}+K$ is on the order of thousands, while $|\mathcal{V}|$ is on the order of tens of thousands. Therefore, the analytical time complexity of RaSeRec is of the same magnitude as the Vanilla and SSL-Augmented SeRec.
4. Experiments
To verify the effectiveness of RaSeRec, we conduct extensive experiments on three benchmark datasets to answer the following research questions (RQs):
  • RQ1: How does RaSeRec perform compared with state-of-the-art SeRec models?
  • RQ2: How much gain can RaSeRec bring to the existing base backbones?
  • RQ3: Can RaSeRec improve long-tailed recommendation?
  • RQ4: Can RaSeRec alleviate preference drift?
  • RQ6: Can RaSeRec perform robustly against the data noise issue?
  • RQ7: How does the performance of RaSeRec vary with different hyper-parameter values? (Appendix F)
  • RQ8: Whether RaSeRec benefits both high-frequency and low-frequency users? (Appendix G)
Table 2. Performance comparison (@5), where the best results are boldfaced and the second-best ones are underlined.
@@ 4. Experiments
Table 3. Performance comparison in terms of different sequential recommendation model backbones.
@@
4.1. Experimental Settings
4.1.1. Datasets
@@
We conduct extensive experiments on three benchmark datasets collected from Amazon (McAuley et al., 2015). We adopt three sub-categories, i.e., Beauty, Sports, and Clothing [^1] http://jmcauley.ucsd.edu/data/amazon/, with different sparsity degrees. Following (Kang and McAuley, 2018; Sun et al., 2019), we discard users and items with fewer than 5 interactions. The statistics of datasets after preprocessing are provided in Appendix B.
4.1.2. Baselines